### v3

# Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources

Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources
Black hole (BH) high mass X-ray binary (HMXB) systems are likely the progenitors of the BH-BH mergers detected in gravitational waves by LIGO/Virgo/KAGRA (LVK). Yet merging BHs reach higher masses () than BHs in HMXBs () and have typically lower spins ( with a larger values tail) than often claimed for BH in HMXBs (). This could suggest that these two classes of systems belong to different populations, but we show here that this does not have to be the case. The difference in masses is easily explained as the known HMXB-BHs are in galaxies with relatively high metallicity, so their progenitor stars are subject to strong mass loss from winds, leading to relatively low mass BH at core collapse. Conversely, LVK can also detect BHs from low-metallicity galaxies that naturally produce more massive stellar-origin BHs. However, a difference in spin is more difficult to explain. Models with efficient angular momentum transport in stellar interiors produces slowly spinning progenitors of both LVK and HMXB BHs. Known HMXBs have orbital periods too long for efficient tidal spin-up, and are also unlikely to have undergone significant accretion spin-up. Instead we show that the derived value of the BH spin depends strongly on how the HMXB accretion disc emission is modelled. We argue that since Cyg X-1 is never observed to be in a soft spectral state, the appropriate spectral models must take into account the Comptonization of the disc photosphere. We show that such models are consistent with low spin values: . This has been confirmed recently by other teams for both Cyg X-1 and LMC X-1, and we show this is also the case for M33 X-7. We conclude that all known HMXB BHs can have low spin, in accordance with the results of stellar evolution models. Hence, current observations are consistent with LVK BHs and HMXB BHs belonging to the same population.
Key Words.:
stars: black holes, compact objects, massive stars†Deceased
1 Introduction
LIGO/Virgo/KAGRA (LVK) interferometers have detected gravitational waves from double black hole (BH-BH) mergers (The LIGO Scientific Collaboration et al. 2023). These merging black holes (BHs) have masses in the range , with many primary (the more massive) BHs having masses of and (Callister & Farr 2024). The most massive LVK event (GW190521) showed two merging BHs with estimated mass of and . The majority of these mergers have low positive effective spin parameters, with the distribution that peaks at
where denotes BH masses, are BH spin magnitudes (: BH angular momentum (AM), : the speed of light and the gravitational constant), and are angles between the system’s angular momentum and BH spins. Out of these mergers show high positive effective spins (mean), while none show high negative spin . However, there are events with moderate negative spin estimates: . Low values of the effective spin parameter may result from small individual BH spin magnitudes (Belczynski et al. 2020b). This finds support in LVK data that shows that BH individual spin magnitudes in BH-BH mergers peak at , although distribution has a long tail; that extends to large values (The LIGO Scientific Collaboration et al. 2023; see Callister & Farr 2024 for slightly different results).
Electromagnetic observations have revealed a population of BHs hosted in binary star systems (e.g., see https://universeathome.pl/universe/blackholes.php and references therein). Out of several known binary configurations hosting BHs, only high-mass X-ray binaries (HMXBs) can potentially lead to BH-BH mergers, as BH companion stars are massive enough to form a BH.There are only three HMXBs known with such massive () companions. These have moderate-mass BHs: LMC X-1 with a star and a BH (Orosz et al. 2009), Cyg X-1 with a star and a BH (Miller-Jones et al. 2021) and M33 X-7 with a star and a BH (Ramachandran et al. 2022 revised down from the previous estimates of a donor and BH in Orosz et al. 2007). (We note that the binary LMC X-3 is not a HMBXB since its stellar companion is less massive () than the BH () Orosz et al. (2014) and won’t form a BH anyway.) Spins of BHs in HMXBs are obtained with two methods namely disc continuum fitting, and reflection spectroscopy. Reflection spectroscopy is complex for soft state spectra (e.g. Tomsick et al. 2018), so here we focus on spin values from disc continuum fitting. These are typically very high111The spins attributed through various methods to LMXB BH vary from very low to very high values, sometimes for the same system (e.g. Reynolds 2021). However, on the one hand these binary systems are not BHBH binary progenitors, on the other, their evolutionary history is still a puzzle (Wiktorowicz et al. 2014).:
Since masses and spins of LVK BHs and HMXB BHs appear to be different, it might seem that these two groups of systems belong to populations with different formation scenarios. Here, we show that all these BHs may instead form a single population. The difference in BH mass arises naturally from the diverse formation environments: all known HMXB-BH are in local group galaxies with relatively high metallicity, so they have strong stellar winds before the core collapse, resulting in lower black hole masses compared to those forming in low metallicity galaxies which are within the sensitivity range of LVK. We demonstrate the difference in expected black hole masses explicitly using a population synthesis model. This model also reproduces the spin values in the LVK BHs, for models where high spins are produced only by tidal spin-ups in extremely close binaries. The known HMXB-BH systems are in binary systems which are not close enough for tidal spin-up. Hence the population synthesis models predict they should not have high spins, which conflicts with the reported high values. However, the disc continuum method relies on the emission thermalizing. If instead the disc is covered by a warm, optically thick skin, so that the emission is Comptonised, the spin values are all consistent with (Belczynski et al. 2021; Zdziarski et al. 2024a, b; this paper). With this interpretation, the LVK BH-BH mergers and the HMXBs can be members of the same binary star population.
2 BH Masses
To demonstrate the unicity of the BH population, we use the population synthesis code StarTrack (Belczynski et al. 2002, 2008) and employ the delayed core-collapse supernova (SN) engine for neutron star/BH mass calculation (Fryer et al. 2012) that does not produce a lower mass gap between NSs and BHs. We use a model in which the pair-instability (upper) mass gap is calculated with non-standard fusion reaction rates and heavy mixing for massive stars that allows for BHs to form up to (Farmer et al. 2020; Costa et al. 2021). Such model was already developed by Belczynski (2020) in the context of the detected at the low metallicity GW190521 LVK merger detected at distance of Gpc Abbott et al. (2020)).
In Figure 1 we show the dependence of the Helium core mass on metallicity for single stars obtained with the stellar wind treatment presented in Belczynski et al. (2010a). The helium core mass is a good approximation of the BH mass since binary interactions (Roche Lobe Overflow or common envelope evolution) tend to remove the H-rich envelope from stars in close binaries (progenitors of BH-BH mergers if these form through isolated binary evolution). The lower the metallicity, the lower the opacity in stellar atmospheres, and the lower the wind mass loss, which produces progressively higher mass BHs with decreasing metallicity.
Cyg X-1 is located in the Milky Way disc that hosts stars with about solar metallicity (, Asplund et al. (2009)). For such metallicity, we obtain a maximum BH mass of , which is consistent with the highest mass stellar-origin BH known in the Milky Way for the same Cyg X-1 ( Miller-Jones et al. (2021)). LMC X-1 is located in the Large Magellanic Cloud () and M33 X-7 in M33 galaxy (). In this metallicity range, BHs are not expected to be more massive than .
The LVK BH-BH mergers have already been detected (O3) to significant redshifts of . The detected BH-BH mergers may have formed even at much larger redshifts due to non-zero delay time between star formation and BH-BH mergers (Dominik et al. 2012; Fishbach & Kalogera 2021). Therefore, galaxies and stars that can produce BH-BH mergers cover a wide range of metallicity. It is even claimed that Population III (metal-free) stars () can contribute to LVK detections (Kinugawa et al. 2014). Assuming a metallicity range for LVK we obtain a broad range of BH mass: . This range is consistent with the LVK BH mass estimates in BH-BH mergers. The recent discovery of a BH in the binary Gaia BH3 (Gaia Collaboration et al. 2024) strongly supports the scenario according to which metal-poor massive stars are progenitors of the high-mass BHs observed by LVK (see, e.g. Olejak et al. 2020).
In conclusion, the same model of stellar evolution produces correct mass ranges for the three discussed HMXB BHs (narrow) and LVK BHs (wide).
3 BH spins: stellar evolution
Population synthesis models are also used to predict the BH spin distribution. In most cases, they produce low-spinning BHs () from non-interacting stars and low- to rapidly-spinning BHs () for Wolf-Rayet (WR) stars that were subject to tidal spin-up in close binaries (Belczynski et al. 2020b; Bavera et al. 2020). Such models can explain the low and high effective spins of LVK BH-BH mergers (Olejak & Belczynski 2021).
The low BH spins obtained by population synthesis models are the consequence of the assumed efficient angular momentum transport from stellar cores to envelopes, an effect that has been now well confirmed by observations (Aerts et al. 2019; Langer 2012) consistent with the action of the magnetic Tayler (1973) instability (Spruit 2002a; Fuller et al. 2019; Eggenberger et al. 2022; Petitdemange et al. 2024). The heuristic Spruit (2002a) version of the mechanism used in population synthesis codes has been recently confirmed by numerical calculations of Petitdemange et al. (2024).
Therefore unless their progenitors in HMXBs belong to a special category of stars with uncoupled cores and envelopes, BHs in these systems should be formed with low spins (). The known HMXB orbital periods are too long for tidal interactions to spin up the stars before they form BHs (the size of the present BH’s companion would not fit into a days orbit of the tidal locking regime; see e.g. Belczynski et al. 2020a) so if their spins are really large they must have have gained angular momentum by accretion. To spin up a BH from to one has roughly to double its mass (Bardeen 1970). For Cyg X-1 this would require accreting in less than Myr (Miller-Jones et al. 2021), i.e an accretion rate , or higher by factor of 100 if the accretion lifetime of Cyg X-1 is only a few tens of thousands of years Miller-Jones et al. 2021). This gives an accretion rate of where is the Eddington accretion rate with an assumed efficiency of . Contrary to the still expressed opinion (see e.g. Miller-Jones et al. 2021), the Eddington rate is not a maximum theoretical limit (Begelman 1979) and black holes can accrete from discs at rates up to at least a few tens of (Kitaki et al. 2021; Yoshioka et al. 2022, but see Hu et al. 2022), so that, an accretion rate in Cyg X-1 is not, from this point of view, extravagant. The problem is to find an evolutionary path that would take a high-mass binary through such a super-Eddington phase and produce the observed Cyg X-1, whose O supergiant donor star has yet to fill its Roche lobe. This is unlikely to happen for accretion from wind, even when it is L1-focused (Wiktorowicz et al. 2021). Therefore, according to the best stellar physics available, BH spins in the known HMXBs should be low, with – consistent with LVK BH spins (Fishbach & Kalogera 2022) and estimates of BH spin in HMXBs by Belczynski et al. (2021); Zdziarski et al. (2024a, b), but clearly incompatible with the generally accepted high spin values in these systems (Miller & Miller 2015). These high spin values have been obtained by assuming the presence of a standard disc in the system (see Sect. 7 for the standard-disc definition), which can be justified only if the system is in a soft spectral state. In the following we will argue that since Cyg X-1 has never been observed in such a state, the high black hole spin from disc continuum fitting is not a firm requirement.
4 BH spins from disc spectral fitting: HMXBs
[…]
Standard accretion disc models balance gravitational heating from viscous torques with blackbody cooling, resulting in an optically thick, geometrically thin structure with effective temperature at each radius, . This gives a total spectrum, which is a sum of all radii of blackbody spectra of different temperatures. The emissivity peaks around , so the peak temperature is K, i.e. is a factor higher for the same disc luminosity at high spin than at low spin.
In all X-ray binaries, one observes cycles during which the brightness and spectral hardness form a characteristic “tortoise-head” pattern. In general, the spectrum contains both thermal disc emission, together with a power law tail extending to much higher energies, clearly showing that there is an additional non-disc component to the accretion flow. Most transients show a phase in this cycle where the spectrum is sufficiently soft to be termed disc dominated, where the assumption of accretion disc assumption is likely to be reliable. The problem is that not all X-ray sources complete the full cycle. As we will show below, this is the case for Cyg X-1.
Indeed, there is clear evidence that in Cyg X-1, the accretion flow has not formed a standard black body disc. The blue points in Figure 2 show the hardness-intensity diagram of the LMXB GX339-4. This source picks out a well-defined track, transitioning from low/hard (lower right) to disc-dominated (upper left, hardness ratio below ), with fast transitions (horizontal) between the two states. The red points compare data from Cyg X-1 (both data taken from Belloni 2010). The mass and distance of GX339-4 are not well known, so we have scaled the relative count rate (brightness) between the two systems to match the lower transition luminosity. We can see how GX339-4 transitions from low/hard state (hardness ratio ; the ratio is defined as the ratio of counts in the energy bands and keV) to high/soft state (hardness ratio ). In contrast, Cyg X-1 never reaches as far to the left as seen in the GX339-4 disc-dominated states.
Hence, Cyg X-1 always has a strong tail of emission to higher energies (e.g. Gierliński et al. 1999; Gou et al. 2009; Walton et al. 2016), making it stuck in the more complex intermediate states where there is no reason to believe that the accretion flow forms a standard accretion disc. Instead, the soft component in intermediate state spectra is often better fit by optically thick, warm (few keV) Comptonization (see, for e.g. transition spectra from the LMXB-BH MAXI J1820+070 in Fig. 2 of Kawamura et al. 2023).
Cyg X-1 has a higher mass BH (; Miller-Jones et al. 2021) than GX339-4 (; Heida et al. 2017), but this should make its disc spectra softer rather than harder since the ISCO is farther out for more massive BHs, making the disc temperature lower. A hotter disc temperature could instead result if Cyg X-1 has a substantially larger spin than GX339-4, but the reflection and reverberation spin estimates for GX339-4 are both extremely high. Therefore, there is no evidence that a standard disc forms in the whole available set of observations of Cyg X- 1, making this system’s standard disc-fitting method of BH spin determination inapplicable.
Instead, the strong soft component need not be emitted by a simple blackbody disc, where the energy is dissipated mainly on the midplane but can have heating further up into the photosphere. This forms a warm layer on the top of the disc, making it appear hotter and giving higher spin when fit with standard disc models. Such a warm layer on the disc is sometimes required in the black hole binaries (e.g. in very high/intermediate type spectra: Kubota et al. 2001; Remillard & McClintock 2006) and is generic in Active Galactic Nuclei (Porquet et al. 2019; Gierliński & Done 2004b; Petrucci et al. 2018). On the other hand it is well known that the simplistic assumption of dissipation concentrated at the disc midplane (usually in the form of dissipation proportional to pressure) is in contradiction with both observations (see e.g., Hubeny & Long 2021) and numerical simulations (Blaes 2014).
Thus, there is no physical reason to assume that a standard accretion disc emits the observed X-ray spectrum of Cyg X-1. In fact, there are good reasons to think that such an assumption is inappropriate for this system. There is no “canonical” model describing the inner disc structure in X-ray binaries – no model correctly describes the “tortoise-head” cycle. Hence, one is allowed to test various options that would reproduce the observed spectrum and check if the deduced spin value is model-dependent, and if it is, which model corresponds best to the value obtained from stellar evolution. A similar approach was adopted by Zdziarski et al. (2024a, b).
5 Cyg X-1 spin from the soft component
We confront two models of Cyg X-1 emission. We use data from the softest state ever seen (observed by Suzaku in 2013; magenta star in Fig. 2), when the soft component is at its most dominant, and so uncertainties from modelling the high energy coronal tail are smallest (observation B4 from Kawano et al. 2017, see this paper for details of the data). We fit these data using the xspec spectral fitting package and all the models below include absorption along the line of sight using tbabs with column density as a free parameter (see Appendix A for the details). We show results from a series of model fits in Figure 3, where the original (absorbed) X-ray Suzaku data is grey, with the reconstructed (corrected for absorption) spectrum shown with orange points. We assume here the standard mass and distance, and fix the disc inclination to the binary inclination of .
We first fit the standard disc model to these data. The results are shown in the left panel of Figure 3. This includes all special and general relativistic ray tracing effects, as well as allowing a correction to the blackbody temperatures to account for the incomplete thermalisation expected from full models of the vertical disc structure (colour temperature correction, fixed to ). We allow a fraction of these disc spectra to be Compton upscattered to form the power law tail, modelled using the standard formalism, which assumes the electron energy is typically higher than keV. We recover the usual, very high black hole spin () seen in all standard-disc fits to the soft component (Zhao et al. 2021a; Zdziarski et al. 2024a, b).
The middle panel of Figure 3 shows what happens when we additionally allow a warm skin on the top of the disc. We model this using another Compton upscattering model, which allows the electron temperature to be a free parameter, as well as the scattered fraction (set by the optical depth of the skin). This is formally a much better fit to the data, even accounting for the additional free parameters: optical depth and temperature of the warm disc skin. A simple F-test gives a probability of the improvement being random as . This is well below the generally accepted limit of for a result to be significant. More importantly, the best fit reduces BH spin to . The BH spin here is extremely poorly constrained as the soft component shape in this model is set more by the electron temperature of the skin than by the temperature of the disc at the ISCO. We illustrate this in the right panel of Figure 3, where we show a fit with almost the same but for , as predicted by standard stellar evolution.
This demonstrates that formal solutions for BH spin from spectra with their reported extremely small error bars or very definitive values (e.g., at level (Zhao et al. 2021b)) are model-dependent and are not the only solution (see also Zdziarski et al. 2024a, b who arrive at the same conclusion).
5.1 Can a disc form in Cyg X-1 for low BH spin?
Modelling of the disc continuum spectra assumes that an accretion disc can form around the BH for any value of the BH spin and/or any orientation of the accretion disc with respect to the spin of the BH. In a binary system, a disc will form only if the circularisation radius of the transferred matter is larger than the radius of the accreting body. In other words, a disc can form if the angular momentum of the accreting matter is large enough to allow it to circle around the accretor. For wind (Bondi-Hoyle; Bondi 1952) accretion the condition is (Frank et al. 2002)
where is the mass of the donor star, the orbital separation, the wind velocity and the escape velocity from the donor. For the parameters of Cyg X-1, when , cm but cm, so that e.g. for no disc can form. Sen et al. (2021) analysed the problem in great detail and showed that for the observed binary and stellar parameters of Cyg X-1 (Miller-Jones et al. 2021), a maximally spinning BH may be required to form an accretion disc in the prograde orientation when the BH accretes from a wind.
However, we note that the O star companion in Cyg X-1 is nearly filling its Roche lobe (Roche-lobe filling factor is greater than 0.9, Miller-Jones et al. 2021), where the formation of a focused accretion stream is expected (Blondin et al. 1991; Hadrava & Čechura 2012; El Mellah et al. 2019). In Cyg X-1, the existence of a focused accretion stream has also been observationally verified (Miller et al. 2005; Poutanen et al. 2008; Hanke et al. 2009). A high Roche lobe filling factor has been found for LMC X-1 (Orosz et al. 2009) and the O donor in M33 X-7 effectively fills its Roche lobe (although not in a standard way, see Ramachandran et al. 2022).
In such a configuration, spherically symmetric Bondi-Hoyle accretion is not accurate. The tidal and gravitational effects of the BH distort the shape of the O star and wind streamlines from the O star, respectively (El Mellah et al. 2019; Hirai & Mandel 2021). Hirai & Mandel (2021) showed that for Cyg X-1, the effects of rotation and gravity darkening can lead to the formation of a prograde accretion disc around a non-rotating BH when the Roche lobe filling factor of the O star is greater than 0.8.
Hence, the criterion of the formation of an accretion disc in Cyg X-1 does not rule out any assumptions on the spin of the BH. The high Roche lobe filling factor of the O star enables the formation of a focused wind with sufficient angular momentum to make a prograde accretion disc. In such case the circularisation radius for the bulk fo the accretion flow is of the order of the solar radius (see, e.g., Frank et al. 2002), orders of magnitude larger than the ISCO radius. The above arguments, together with our modelling of the disc continuum spectra imply that the observations of Cyg X-1 are consistent with the statement that its BH can also be slowly spinning.
6 M33 X-7 BH spin from the soft component
The source X-7 in the nearby galaxy M33 is an eclipsing HMBH on a quasi-circular, 3.45 day orbit. In this system, according to phase-resolved simultaneous HST and XMM-Newton observations by Ramachandran et al. (2022), an BH accretes matter lost from its O9II, metal-poor, supergiant companion. The previous mass estimate by Orosz et al. (2007) had a higher BH mass of (and companion).
Assuming the older BH mass, Liu et al. (2008) used the continuum fitting method to get a BH spin . We have used the standard continuum method to determine the M33 X-7 spin value for the new BH mass and obtained (see Appendix B), as expected. With this new spin value, the M33 X-7 BH is no longer in the very-rapidly rotating class, to which Cyg X-1 with and LMC X-1 with belong. Nevertheless, since we have concluded above that the spin determination in Cyg X-1 is strongly model-dependent and Zdziarski et al. (2023, 2024b) arrived to same conclusion regarding LMC X-1, we thought it useful to check if the same applies to M33 X-7. Similarly to these sources, when a warm-skin Comptonizing model is used, the black hole spin is completely unconstrained, with acceptable values ranging from -1 to at confidence level (see Appendix B).
Below, we will discuss the possible reason for this common property of the only three HMBXs that are potential progenitors of BHBH binary systems.
7 Discussion
We have shown that the X-ray spectra of Cyg X-1 and M88 X-7 can be fitted with both standard disk and warm Comptonised disc models (see also Zdziarski et al. 2024a, b for the same conclusion about Cyg X-1 and LMC X-1). We have also noticed that even in its softest state, the spectrum of Cyg X-1 is not soft enough to justify a standard-disc model fit, so there are very good reasons to consider alternatives.
In order to avoid ambiguities and misunderstandings, one has to define precisely what is understood by fitting spectra with a “standard disc” model. This model is describing an axially symmetric, stationary, geometrically thin Keplerian accretion flow. It is assumed that the angular momentum transport is provided by a “viscous” (in astrophysical discs, the viscosity must be anomalous) torque and that the local dissipation of the same torque is the source of the energy radiated away from the disc surface. Then, from mass and angular momentum conservation equations for the vertically averaged disc parameters, assuming thermal equilibrium, one obtains
where is the effective temperature and contains information about the boundary conditions. This is the equation that provides the flux, or effective temperature profile for a Keplerian stationary disc. It does not say that the disc emits radiation like a black body. It contains no information about the vertical energy transport in the disc and is independent of the angular momentum and dissipation mechanism it only assumes the existence of such mechanism (it is independent of the value of the “viscosity parameter , it does not even assume the existence of such a parameter). Eq. (3) is the basis of the “standard” disc X-ray spectrum model. There, one assumes that the colour temperature is close to the effective temperature. i.e., that the disc, locally radiates approximately like a black body. Since the accretor is a BH, one adds to the model general-relativistic effects, including radiation back-falling onto the disc, and appropriate inner boundary conditions (usually zero torque at ISCO). Since the observed spectra are not strictly a sum of black-bodies, a disc annulus emission is represented by a colour-corrected (or diluted) black-body
where is the intensity, is the Planck spectrum, the Planck constant, the radiation frequency and is a colour temperature correction factor representing the extent to which electron scattering through the disc vertical structure is expected to affect the emergent spectrum. This can be calculated by solving the disc vertical radiative transfer equations using a code such as TLUSTY (see Davis & El-Abd 2019; Davis et al. 2005 and references therein). However, contrary to to stellar atmospheres, accretion discs are not in radiative equilibrium, i.e. the divergence of the radiative flux is not equal to zero, but to the (viscous) heating rate per unit volume. Hence the calculations requires the viscous–dissipation vertical structure as input. In the -model of Shakura & Sunyaev (1973), one assumes that this dissipation profile is identical to that of the total pressure, but this is radiation pressure unstable at the observed disc luminosities. Instead, for the disc calculations, the dissipation is assumed to follow the density profile in TLUSTY Done & Davis (2008). Both these prescriptions have maximal dissipation at the mid-plane, whereas MRI simulations often show maximum dissipation closer to the disc surface (see, e.g., Blaes 2014). Some versions of TLUSTY (Hubeny & Hubeny 1998) use a step-wise power-law dependence of viscosity that allows a chosen fraction of the dissipation to occur in the surface layers. Such dissipation profile is needed to reproduce disc spectra of cataclysmic variables Hubeny & Long (2021) and when applied to BH X-ray emitting disc, such dissipation stratification produces disc coronas (Davis & El-Abd 2019) as suggested by Svensson & Zdziarski (1994).
What is, or rather what are the real, physical dissipation profiles is still unknown since numerical simulations cannot yet produce realistic vertical structures of geometrically thin accretion discs. That is why, disc models with warm surface layers are as physically legitimate as the presumed “standard” model, especially when there is strong evidence as in Cyg X-1 that the structure is not a standard disc.
One should also point out that the -disc model cannot explain the brightness–hardness ratio tortoise-head hysteresis observed in transient X-ray sources. The physical reason for this variability is still unknown. Observations (e.g. Done et al. 2007; Basak & Zdziarski 2016; Basak et al. 2017) and models (e.g., Dubus et al. 2001) suggests that accretion discs in BH X-ray binaries are truncated, but the physical mechanism of this excision is still a mystery.
It does not mean that the -disc model on which the “standard” spectral model is based, is to be rejected. This model has found many successful applications. The soft state spectra in transients show peak disk temperature and total luminosity changing together with , indicating a constant inner radius for the disc and constant Done et al. (2007). Even the radial dependence of the disc temperature shows the predicted behaviour in eclipsing cataclysmic variables in their disk (outbursting) states (Horne & Cook 1985), as well as showing the very different radial profile expected from non-equilibrium discs in quiescent dwarf-novae; Lasota 2001.
None of the known HMBH show a convincing soft state relation which characterises a standard disc (see Sect. 4). Cyg X-1 is never in a soft state, while in LMC X-1, Gierliński & Done (2004a) discovered that the disc luminosity follows instead a track. Zdziarski et al. (2024a) found that such a anti-correlation between luminosity and temperature can be naturally explained as an effect of the warm Comptonization layer on top of a standard disc. The luminosity variations in M88 X-7 are not large enough to determine whether a relation is present 222Liu et al. (2008) present a variability in counts (not in counts per second) which reflects mostly the variability of the exposure time..
The three BHHMBXs, Cyg X-1, LMC X-1 and M33 X-7, whose spectra are well fitted by a model of a disc covered with warm comptonising layer have one feature in common: the accrete not only from focused stream but also capture some matter from the strong wind of the BH companion. It is natural to speculate that this is the source of the warm layer, making it a ubiquitous feature of the HMBH spectra.
8 Conclusion
The difference between HMXB and LVK BH masses has already been noticed after the first gravitational wave detection and attributed to metallicity effect (Abbott et al. 2016) as anticipated by Belczynski et al. (2010b). Taking into account that interferometric detectors of gravitational waves sample a volume that is a few hundred thousand times larger than the X-ray telescopes (for HMXB BH mass determination), the difference in BH masses between LVK and HMXBs is not surprising as it implies a huge difference in the metallicity range sampled (see Fig. 1).
While assessing the problem of BH mass difference between HMXB and gravitational–wave sources is rather straightforward, addressing the apparent conflicting spin values is more complicated. At face value, the spin values attributed to the BH in the HMXBs that are massive enough to form LVK sources (Cyg X- 1, LMC X-1, M33 X-7) seem to require two distinct populations: slowly spinning (LVK) and rapidly spinning (HMXBs) BHs.
Since the majority of LVK BHs are slowly spinning (), we argue, under the assumption that they are stellar-origin BHs (i.e., are not primordial BHs: Hawking 1971, but see Mroz et al. 2024), that it must indicate that massive stars are subject to efficient angular momentum transport (Spruit 2002b; Fuller et al. 2019; Eggenberger et al. 2022; Petitdemange et al. 2024). This leads to good agreement of predicted BH spins in BH-BH mergers with LVK data (Belczynski et al. 2020b; Bavera et al. 2020). These slowly spinning BHs can form in isolated binary evolution and various dense (open, globular, nuclear) cluster environments. A small fraction of LVK BHs may be moderately or even rapidly spinning (see the long tail of BH spins with : left panel of Fig.15 in The LIGO Scientific Collaboration et al. 2023). These moderately/rapidly spinning BHs can also be explained in the framework of efficient angular momentum transport. Although the star rotation slows down during its evolution towards the formation of a BH, the tidal spin-up of a WR star (BH progenitor) in close BH+WR binary systems (BH-BH progenitors) can easily result in a moderately or rapidly spinning BH (Olejak & Belczynski 2021). In dense clusters, dynamical interactions may lead to the formation of BH-BH mergers with components that are second or even third-generation BHs (formed by earlier BH-BH mergers) that naturally have significant spins (Rodriguez et al. 2019).
Under the natural assumption that BHs in HMXBs are formed directly from massive stars and within the paradigm of efficient angular momentum transport, these BHs also must have low spins (; e.g., Belczynski et al. 2020b). We have demonstrated that this conclusion is fully consistent with BH spin estimate based on fitting the disc spectra in HMXBs with models in which the disc is covered by a warm Comptonized layer. In particular, we show that with this model, the X-ray spectra of the softest, but still pretty hard, state of Cyg X-1 is consistent with a slowly spinning BH as confirmed by Zdziarski et al. (2024b) for another set of data. The same is true of X-7 in M33 whose BH spin is completely unconstrained when its soft X-ray spectra are fitted with the warm–skin model. The same warm–skin solutions also give a low spin for LMC X-1 (Zdziarski et al. 2024a. We have shown that the warm–skin solution is as physically plausible as the so–called standard solution, especially in the case of systems that are not in a generic soft X-ray spectral state. We speculate that this warm skin is ubiquitous in the HMXB-BH due to the presence of additional accretion from lower angular momentum material in the stellar wind as well as Roche Lobe Overflow.
Thus, we conclude that the LVK and HMXB BHs are consistent with having the same origin, that is, both classes of BHs form directly from stars with efficient angular momentum transport.
Acknowledgements.
Unfortunately, Chris Belczyński died while working on the re-submission of the present paper. We will miss his deep knowledge, inexhaustible enthusiasm, and kindness. We thank Jerome Orosz, Henric Krawczynski, Andrew King, Aleksandra Olejak and Andrzej Zdziarski for very useful comments on this study. This research has made use of data obtained from the Chandra Data Archive and the Chandra Source Catalog. We are grateful to the anonymous referee for very helpful suggestions and comments. KB was supported by the Polish National Science Center (NCN) grant Maestro (2018/30/A/ST9/00050). CD acknowledges support from STFC through grant ST/P000541/1. SH acknowledges support from STFC through the studentship grant ST/V506643/1. KS is funded by the National Science Center (NCN), Poland, under grant number OPUS 2021/41/B/ST9/00757.References
- Abbott et al. (2016) Abbott, B. P., Abbott, R., Abbott, T. D., et al. 2016, ApJ, 818, L22
- Abbott et al. (2020) Abbott, R., Abbott, T. D., Abraham, S., et al. 2020, Phys. Rev. Lett., 125, 101102
- Abramowicz & Lasota (1980) Abramowicz, M. A. & Lasota, J. P. 1980, Acta Astron., 30, 35
- Aerts et al. (2019) Aerts, C., Mathis, S., & Rogers, T. M. 2019, ARA&A, 57, 35
- Arnaud (1996) Arnaud, K. A. 1996, in Astronomical Society of the Pacific Conference Series, Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes, 17
- Asplund et al. (2009) Asplund, M., Grevesse, N., Sauval, A. J., & Scott, P. 2009, ARA&A, 47, 481
- Bardeen (1970) Bardeen, J. M. 1970, Nature, 226, 64
- Basak & Zdziarski (2016) Basak, R. & Zdziarski, A. A. 2016, MNRAS, 458, 2199
- Basak et al. (2017) Basak, R., Zdziarski, A. A., Parker, M., & Islam, N. 2017, MNRAS, 472, 4220
- Bavera et al. (2020) Bavera, S. S., Fragos, T., Qin, Y., et al. 2020, A&A, 635, A97
- Begelman (1979) Begelman, M. C. 1979, MNRAS, 187, 237
- Belczynski (2020) Belczynski, K. 2020, ApJ, 905, L15
- Belczynski et al. (2010a) Belczynski, K., Bulik, T., Fryer, C. L., et al. 2010a, ApJ, 714, 1217
- Belczynski et al. (2010b) Belczynski, K., Dominik, M., Bulik, T., et al. 2010b, ApJ, 715, L138
- Belczynski et al. (2021) Belczynski, K., Done, C., Hagen, S., Lasota, J. P., & Sen, K. 2021, arXiv e-prints, arXiv:2111.09401v2
- Belczynski et al. (2020a) Belczynski, K., Hirschi, R., Kaiser, E. A., et al. 2020a, ApJ, 890, 113
- Belczynski et al. (2002) Belczynski, K., Kalogera, V., & Bulik, T. 2002, ApJ, 572, 407
- Belczynski et al. (2008) Belczynski, K., Kalogera, V., Rasio, F. A., et al. 2008, ApJS, 174, 223
- Belczynski et al. (2020b) Belczynski, K., Klencki, J., Fields, C. E., et al. 2020b, A&A, 636, A104
- Belloni (2010) Belloni, T. M. 2010, States and Transitions in Black Hole Binaries, ed. T. Belloni, Vol. 794, 53
- Blaes (2014) Blaes, O. 2014, Space Sci. Rev., 183, 21
- Blondin et al. (1991) Blondin, J. M., Stevens, I. R., & Kallman, T. R. 1991, ApJ, 371, 684
- Bondi (1952) Bondi, H. 1952, MNRAS, 112, 195
- Callister & Farr (2024) Callister, T. A. & Farr, W. M. 2024, Physical Review X, 14, 021005
- Costa et al. (2021) Costa, G., Bressan, A., Mapelli, M., et al. 2021, MNRAS, 501, 4514
- Davis et al. (2005) Davis, S. W., Blaes, O. M., Hubeny, I., & Turner, N. J. 2005, ApJ, 621, 372
- Davis & El-Abd (2019) Davis, S. W. & El-Abd, S. 2019, ApJ, 874, 23
- Dominik et al. (2012) Dominik, M., Belczynski, K., Fryer, C., et al. 2012, ApJ, 759, 52
- Done & Davis (2008) Done, C. & Davis, S. W. 2008, ApJ, 683, 389
- Done et al. (2007) Done, C., Gierliński, M., & Kubota, A. 2007, A&A Rev., 15, 1
- Dubus et al. (2001) Dubus, G., Hameury, J. M., & Lasota, J. P. 2001, A&A, 373, 251
- Eggenberger et al. (2022) Eggenberger, P., Moyano, F. D., & den Hartogh, J. W. 2022, A&A, 664, L16
- El Mellah et al. (2019) El Mellah, I., Sander, A. A. C., Sundqvist, J. O., & Keppens, R. 2019, A&A, 622, A189
- Farmer et al. (2020) Farmer, R., Renzo, M., de Mink, S., Fishbach, M., & Justham, S. 2020, arXiv e-prints, arXiv:2006.06678
- Fishbach & Kalogera (2021) Fishbach, M. & Kalogera, V. 2021, ApJ, 914, L30
- Fishbach & Kalogera (2022) Fishbach, M. & Kalogera, V. 2022, ApJ, 929, L26
- Frank et al. (2002) Frank, J., King, A., & Raine, D. J. 2002, Accretion Power in Astrophysics: Third Edition (CUP)
- Fryer et al. (2012) Fryer, C. L., Belczynski, K., Wiktorowicz, G., et al. 2012, ApJ, 749, 91
- Fuller et al. (2019) Fuller, J., Piro, A. L., & Jermyn, A. S. 2019, MNRAS[arXiv:1902.08227]
- Gaia Collaboration et al. (2024) Gaia Collaboration, Panuzzo, P., Mazeh, T., et al. 2024, arXiv e-prints, arXiv:2404.10486
- Gierliński & Done (2004a) Gierliński, M. & Done, C. 2004a, MNRAS, 347, 885
- Gierliński & Done (2004b) Gierliński, M. & Done, C. 2004b, MNRAS, 349, L7
- Gierliński et al. (1999) Gierliński, M., Zdziarski, A. A., Poutanen, J., et al. 1999, MNRAS, 309, 496
- Gou et al. (2009) Gou, L., McClintock, J. E., Liu, J., et al. 2009, ApJ, 701, 1076
- Gou et al. (2014) Gou, L., McClintock, J. E., Remillard, R. A., et al. 2014, ApJ, 790, 29
- Hadrava & Čechura (2012) Hadrava, P. & Čechura, J. 2012, A&A, 542, A42
- Hanke et al. (2009) Hanke, M., Wilms, J., Nowak, M. A., et al. 2009, ApJ, 690, 330
- Hawking (1971) Hawking, S. 1971, MNRAS, 152, 75
- Heida et al. (2017) Heida, M., Jonker, P. G., Torres, M. A. P., & Chiavassa, A. 2017, ApJ, 846, 132
- Hirai & Mandel (2021) Hirai, R. & Mandel, I. 2021, PASA, 38, e056
- Horne & Cook (1985) Horne, K. & Cook, M. C. 1985, MNRAS, 214, 307
- Hu et al. (2022) Hu, H., Inayoshi, K., Haiman, Z., et al. 2022, ApJ, 935, 140
- Hubeny & Hubeny (1998) Hubeny, I. & Hubeny, V. 1998, ApJ, 505, 558
- Hubeny & Long (2021) Hubeny, I. & Long, K. S. 2021, MNRAS, 503, 5534
- Kawamura et al. (2023) Kawamura, T., Done, C., Axelsson, M., & Takahashi, T. 2023, MNRAS, 519, 4434
- Kawano et al. (2017) Kawano, T., Done, C., Yamada, S., et al. 2017, PASJ, 69, 36
- Kinugawa et al. (2014) Kinugawa, T., Inayoshi, K., Hotokezaka, K., Nakauchi, D., & Nakamura, T. 2014, MNRAS, 442, 2963
- Kitaki et al. (2021) Kitaki, T., Mineshige, S., Ohsuga, K., & Kawashima, T. 2021, PASJ, 73, 450
- Kubota et al. (2001) Kubota, A., Makishima, K., & Ebisawa, K. 2001, ApJ, 560, L147
- Langer (2012) Langer, N. 2012, ARA&A, 50, 107
- Lasota (2001) Lasota, J.-P. 2001, New Astronomy Reviews, 45, 449
- Li et al. (2005) Li, L.-X., Zimmerman, E. R., Narayan, R., & McClintock, J. E. 2005, ApJS, 157, 335
- Liu et al. (2008) Liu, J., McClintock, J. E., Narayan, R., Davis, S. W., & Orosz, J. A. 2008, ApJ, 679, L37
- Liu et al. (2010) Liu, J., McClintock, J. E., Narayan, R., Davis, S. W., & Orosz, J. A. 2010, ApJ, 719, L109
- Miller et al. (2005) Miller, J. M., Wojdowski, P., Schulz, N. S., et al. 2005, ApJ, 620, 398
- Miller & Miller (2015) Miller, M. C. & Miller, J. M. 2015, Phys. Rep, 548, 1
- Miller-Jones et al. (2021) Miller-Jones, J. C. A., Bahramian, A., Orosz, J. A., et al. 2021, Science, 371, 1046
- Mroz et al. (2024) Mroz, P., Udalski, A., Szymanski, M. K., et al. 2024, arXiv e-prints, arXiv:2403.02386
- Novikov & Thorne (1973) Novikov, I. D. & Thorne, K. S. 1973, in Black Holes (Les Astres Occlus), 343–450
- Olejak & Belczynski (2021) Olejak, A. & Belczynski, K. 2021, ApJ, 921, L2
- Olejak et al. (2020) Olejak, A., Belczynski, K., Bulik, T., & Sobolewska, M. 2020, A&A, 638, A94
- Orosz et al. (2007) Orosz, J. A., McClintock, J. E., Narayan, R., et al. 2007, Nature, 449, 872
- Orosz et al. (2009) Orosz, J. A., Steeghs, D., McClintock, J. E., et al. 2009, ApJ, 697, 573
- Orosz et al. (2014) Orosz, J. A., Steiner, J. F., McClintock, J. E., et al. 2014, ApJ, 794, 154
- Petitdemange et al. (2024) Petitdemange, L., Marcotte, F., Gissinger, C., & Daniel, F. 2024, A&A, 681, A75
- Petrucci et al. (2018) Petrucci, P. O., Ursini, F., De Rosa, A., et al. 2018, A&A, 611, A59
- Porquet et al. (2019) Porquet, D., Done, C., Reeves, J. N., et al. 2019, A&A, 623, A11
- Poutanen et al. (2008) Poutanen, J., Zdziarski, A. A., & Ibragimov, A. 2008, MNRAS, 389, 1427
- Ramachandran et al. (2022) Ramachandran, V., Oskinova, L. M., Hamann, W. R., et al. 2022, A&A, 667, A77
- Remillard & McClintock (2006) Remillard, R. A. & McClintock, J. E. 2006, ARA&A, 44, 49
- Reynolds (2021) Reynolds, C. S. 2021, ARA&A, 59, 117
- Rodriguez et al. (2019) Rodriguez, C. L., Zevin, M., Amaro-Seoane, P., et al. 2019, Phys. Rev. D, 100, 043027
- Sen et al. (2021) Sen, K., Xu, X. T., Langer, N., et al. 2021, A&A, 652, A138
- Shakura & Sunyaev (1973) Shakura, N. I. & Sunyaev, R. A. 1973, A&A, 500, 33
[…]
- Spruit (2002a) Spruit, H. C. 2002a, A&A, 381, 923
- Spruit (2002b) Spruit, H. C. 2002b, A&A, 381, 923
- Steiner et al. (2009) Steiner, J. F., Narayan, R., McClintock, J. E., & Ebisawa, K. 2009, PASP, 121, 1279
- Svensson & Zdziarski (1994) Svensson, R. & Zdziarski, A. A. 1994, ApJ, 436, 599
- Tayler (1973) Tayler, R. J. 1973, MNRAS, 161, 365
- The LIGO Scientific Collaboration et al. (2023) The LIGO Scientific Collaboration, the Virgo Collaboration, the KAGRA Collaboration, et al. 2023, Physical Review X, 13, 011048
- Thorne (1974) Thorne, K. S. 1974, ApJ, 191, 507
- Tomsick et al. (2018) Tomsick, J. A., Parker, M. L., García, J. A., et al. 2018, ApJ, 855, 3
- Tripathi et al. (2020) Tripathi, A., Zhou, M., Abdikamalov, A. B., et al. 2020, ApJ, 897, 84
- Walton et al. (2016) Walton, D. J., Tomsick, J. A., Madsen, K. K., et al. 2016, ApJ, 826, 87
- Wiktorowicz et al. (2014) Wiktorowicz, G., Belczynski, K., & Maccarone, T. 2014, in Binary Systems, their Evolution and Environments, ed. R. de Grijs, 37
- Wiktorowicz et al. (2021) Wiktorowicz, G., Lasota, J.-P., Belczynski, K., et al. 2021, ApJ, 918, 60
- Yoshioka et al. (2022) Yoshioka, S., Mineshige, S., Ohsuga, K., Kawashima, T., & Kitaki, T. 2022, PASJ, 74, 1378
- Zdziarski et al. (2024a) Zdziarski, A. A., Banerjee, S., Chand, S., et al. 2024a, ApJ, 962, 101
- Zdziarski et al. (2024b) Zdziarski, A. A., Chand, S., Banerjee, S., et al. 2024b, ApJ, 967, L9
- Zdziarski et al. (1996) Zdziarski, A. A., Johnson, W. N., & Magdziarz, P. 1996, MNRAS, 283, 193
- Zdziarski et al. (2020) Zdziarski, A. A., Szanecki, M., Poutanen, J., Gierliński, M., & Biernacki, P. 2020, MNRAS, 492, 5234
- Zdziarski et al. (2023) Zdziarski, A. A., Veledina, A., Szanecki, M., et al. 2023, ApJ, 951, L45
- Zhao et al. (2021a) Zhao, X., Gou, L., Dong, Y., et al. 2021a, ApJ, 916, 108
- Zhao et al. (2021b) Zhao, X., Gou, L., Dong, Y., et al. 2021b, ApJ, 908, 117
Appendix A Fitting Cyg X-1 spectra
We fit Cyg X-1 using two models: one considering a standard disc, and another where the disc is covered by some warm optically thick material (model a and b respectively). In both cases we model the spectra using xspec v.12.13.0 (Arnaud 1996).
For the standard disc we use the xspec model kerrbb (Li et al. 2005). This considers multi-colour black-body emission originating from a disc structure, assuming Novikov & Thorne (1973) emissivity. All relativistic effects are taken into account, including ray-tracing from the disc to the observer. Additionally, this model always assumes that the disc extends to the ISCO.
However, Cyg X-1 always shows a high energy tail that cannot be disc emission. This is generally understood as originating from a hot optically thin inner corona, which Compton up-scatters incident disc photons to higher energies. To model this we convolve our disc emission, kerrbb, with simpl (Steiner et al. 2009). This takes an input seed spectrum (in this case kerrbb) and scatters a fraction of the seed photons into a power-law component (hence emulating a Comptonized spectrum). In xspec syntax, the model is simpl*kerrbb.
For model b we also have the addition of a warm optically thick skin covering the disc. This will also Compton up-scatter the disc photon to higher energies, but not sufficiently so to give a high energy tail. Instead this will give a spectrum that looks similar to a disc, but with the peak shifted to slightly higher energies and a shallower drop-off than the usual Wien tail. We model this using the convolution model thcomp (Zdziarski et al. 1996, 2020). This takes an input seed spectrum and Compton scatters to higher energies, similar to simpl, however unlike simpl thcomp takes the electron temperature of the Comptonizing plasma as an input argument, setting the high energy turn over. Additionally, for this optically thick skin we assume that all input seed photons are Compton scattered, unlike the case for the high-energy tail where only a fraction are scattered. In xspec syntax the model is now csimpl*thcomp*kerrbb.
The line of sight absorption to Cyg X-1 is complex and variable. Hence we need to include an absorption component in both our models, for which we use tbabs. Due to the variable nature of the absorber in Cyg X-1 we leave the column-density in tbabs as a free parameter throughout. We note that in Fig. 3 we show the spectrum corrected for absorption in colour, and the original uncorrected spectrum in grey.
As the data originate from two different instruments, XIS (soft-xrays) and HXD (hard-xrays), though both onboard Suzaku, we incldue a cross-calibration constant in all our models. Throughout this is fixed to 1.15, i.e increasing the model flux in the hard band by a factor 1.15. Finally, the total xspec models are: const*tbabs*csimpl*kerrbb for model a, and const*tbabs*csimpl*thcomp*kerrbb for model b.
Appendix B Re-Visiting M33-X7
Here we re-visit the black hole spin estimate of M33-X7, performed by Liu et al. (2008), using the same methodology as for Cyg X-1. We extract the four Chandra ACIS spectra, referred to as the ’gold’ spectra in Liu et al. (2008). Initial inspection of the data show strongly disc dominated spectra, with no clear non-thermal high energy tail within the observed energy range. Hence we use a slightly simpler model than with Cyg X-1, modelling the continuum as either a pure disc (using kerrbb, Li et al. (2005)) or a disc with a warm Comptonising medium above it (using thcomp*kerrbb, Zdziarski et al. 1996, 2020). We also include the effects of absorption along our line of sight, which will affect the low energy emission, using tbabs. Hence the total xspec models are: tbabs*kerrbb for the pure disc case, and tbabs*thcomp*kerrbb for the warm Comptonised case.
We note that since the analysis of Liu et al. (2008) the black hole mass of M33 X-7 has been updated from (Orosz et al. 2007) to (Ramachandran et al. 2022). The main effect of this will be to lower the black hole spin estimate in the case of a pure disc, as a lower black hole mass will shift the peak of the spectrum to higher energies, much in the same way as an increase in black hole spin will do. For completeness we fit the data using both the new and previous mass estimates, to highlight that our conclusions are not sensitive to this change.
Fig. 4 shows the four ACIS spectra along with their respective model fits, for the pure disc (left) and warm Comptonised medium (right). The data show a change in normalisation between some of the spectra, corresponding to a change in the mass-accretion rate (see e.g the difference between the green/black data/model and the magenta/orange data/model in Fig. 4). Hence, during the spectral fitting we allow the mass-accretion rate to vary between the datasets, but then tie all other parameters as these should not vary between observations. The spectra were grouped such that each energy bin contained a minimum of 20 counts, allowing for the use of statistics.
Fig. 5 shows the change in as we step the fit through the possible values for black hole spin, with a standard disc on the left and a disc covered by a warm Comptonising material on the right. The standard disc clearly gives a strongly constrained value for the black hole spin, and in fact for the old mass estimate we recover the value derived in Liu et al. (2008) of . However, it is clear that a change in mass will affect this result, with the new mass estimate reducing the black hole spin to . However, the accretion flow is not necessarily characterised by a standard disc. Instead it is likely that it could be covered in a warm Comptonising material, as discussed previously for Cyg X-1 and in LMC X-1 (Zdziarski et al. 2024a, b). Unlike the case of a pure disc, this gives completely unconstrained black hole spin, with acceptable values ranging from -1 to at confidence. As with both Cyg X-1 and LMC X-1, this demonstrates the strongly model dependent nature of black hole spin estimates.

### v4

@@
-Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources
-Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources
-Black hole (BH) high mass X-ray binary (HMXB) systems are likely the progenitors of the BH-BH mergers detected in gravitational waves by LIGO/Virgo/KAGRA (LVK).
-Yet merging BHs reach higher masses () than BHs in HMXBs () and have typically lower spins ( with a larger values tail) than often claimed for BH in HMXBs ().
-This could suggest that these two classes of systems belong to different populations, but we show here that this does not have to be the case.
-The difference in masses is easily explained as the known HMXB-BHs are in galaxies with relatively high metallicity, so their progenitor stars are subject to strong mass loss from winds, leading to relatively low mass BH at core collapse.
-Conversely, LVK can also detect BHs from low-metallicity galaxies that naturally produce more massive stellar-origin BHs.
-However, a difference in spin is more difficult to explain.
-Models with efficient angular momentum transport in stellar interiors produces slowly spinning progenitors of both LVK and HMXB BHs.
-Known HMXBs have orbital periods too long for efficient tidal spin-up, and are also unlikely to have undergone significant accretion spin-up.
-Instead we show that the derived value of the BH spin depends strongly on how the HMXB accretion disc emission is modelled.
-We argue that since Cyg X-1 is never observed to be in a soft spectral state, the appropriate spectral models must take into account the Comptonization of the disc photosphere.
-We show that such models are consistent with low spin values: .
-This has been confirmed recently by other teams for both Cyg X-1 and LMC X-1, and we show this is also the case for M33 X-7.
-We conclude that all known HMXB BHs can have low spin, in accordance with the results of stellar evolution models.
-Hence, current observations are consistent with LVK BHs and HMXB BHs belonging to the same population.
+Common origin for black holes in both high mass X-ray binaries and gravitational-wave sources
+Common origin for black holes in both high mass X-ray binaries and gravitational-wave sources
+Black-hole (BH) high-mass X-ray binary (HMXB) systems are likely to be the progenitors of BH-BH mergers detected in gravitational waves by LIGO/Virgo/KAGRA (LVK).
+Yet merging BHs reach higher masses () than BHs in HMXBs () and typically exhibit lower spins ( with a larger values tail) than what is often claimed for BHs in HMXBs ().
+This could suggest that these two classes of systems belong to different populations, but here we show that this may not necessarily be the case.
+The difference in masses is easily explained as the known HMXB-BHs are in galaxies with relatively high metallicity, so their progenitor stars are subject to strong mass loss from winds, leading to relatively low-mass BH at core collapse.
+Conversely, LVK is also able to detect BHs from low-metallicity galaxies that are known to naturally produce more massive stellar-origin BHs.
+However, the difference in spin is more difficult to explain.
+Models with efficient angular momentum transport in stellar interiors produce slowly spinning progenitors for both LVK and HMXB BHs.
+Known HMXBs have orbital periods that are too long for efficient tidal spin-up and are also unlikely to have undergone significant accretion spin-up.
+Instead, we show that the derived value of the BH spin depends strongly on how the HMXB accretion disc emission is modelled.
+We argue that since Cyg X-1 is never observed to be in a soft spectral state, the appropriate spectral models must take into account the Comptonisation of the disc photosphere.
+We show that such models are consistent with low spin values, namely: .
+This was recently confirmed by other teams for both Cyg X-1 and LMC X-1 and here we show this is also the case for M33 X-7.
+We conclude that all known HMXB BHs can exhibit a low spin, in accordance with the results of stellar evolution models.
+Hence, the observations presented in this work are consistent with the LVK BHs and HMXB BHs belonging to the same population.
 Key Words.:
@@
 The most massive LVK event (GW190521) showed two merging BHs with estimated mass of and .
-The majority of these mergers have low positive effective spin parameters, with the distribution that peaks at
-where denotes BH masses, are BH spin magnitudes (: BH angular momentum (AM), : the speed of light and the gravitational constant), and are angles between the system’s angular momentum and BH spins.
-Out of these mergers show high positive effective spins (mean), while none show high negative spin .
-However, there are events with moderate negative spin estimates: .
+The majority of these mergers have low positive effective spin parameters, with the distribution peaking at
+where denotes BH masses, are the BH spin magnitudes (: BH angular momentum, with as the speed of light and as the gravitational constant), while refers to the angles between the system’s angular momentum and BH spins.
+Out of these mergers, of them show high positive effective spins (mean), while none show a high negative spin .
+However, there are three events with moderate negative spin estimates: .
 Low values of the effective spin parameter may result from small individual BH spin magnitudes (Belczynski et al.
 2020b).
-This finds support in LVK data that shows that BH individual spin magnitudes in BH-BH mergers peak at , although distribution has a long tail; that extends to large values (The LIGO Scientific Collaboration et al.
+This is supported by LVK data, which shows that BH individual spin magnitudes in BH-BH mergers peak at ; however, the distribution has a long tail that extends to large values (The LIGO Scientific Collaboration et al.
 2023; see Callister & Farr 2024 for slightly different results).
-Electromagnetic observations have revealed a population of BHs hosted in binary star systems (e.g., see https://universeathome.pl/universe/blackholes.php and references therein).
-Out of several known binary configurations hosting BHs, only high-mass X-ray binaries (HMXBs) can potentially lead to BH-BH mergers, as BH companion stars are massive enough to form a BH.There are only three HMXBs known with such massive () companions.
-These have moderate-mass BHs: LMC X-1 with a star and a BH (Orosz et al.
-2009), Cyg X-1 with a star and a BH (Miller-Jones et al.
-2021) and M33 X-7 with a star and a BH (Ramachandran et al.
-2022 revised down from the previous estimates of a donor and BH in Orosz et al.
+Electromagnetic observations have revealed a population of BHs hosted in binary star systems111e.g.
+see https://universeathome.pl/universe/blackholes.php and references therein.
+Out of several known binary configurations hosting BHs, only high-mass X-ray binaries (HMXBs) can potentially lead to BH-BH mergers, as BH companion stars are massive enough to form a BH.
+There are only three HMXBs known to have such massive () companions and they have moderate-mass BHs: LMC X-1 with a star and a BH (Orosz et al.
+2009); Cyg X-1 with a star and a BH (Miller-Jones et al.
+2021); and M33 X-7 with a star and a BH (Ramachandran et al.
+2022; revised down from the previous estimates of a donor and BH in Orosz et al.
 2007).
-(We note that the binary LMC X-3 is not a HMBXB since its stellar companion is less massive () than the BH () Orosz et al.
-(2014) and won’t form a BH anyway.) Spins of BHs in HMXBs are obtained with two methods namely disc continuum fitting, and reflection spectroscopy.
+We note that the binary LMC X-3 is not a HMBXB since its stellar companion is less massive () than the BH () Orosz et al.
+(2014) and will not form a BH anyway.
+The spins of BHs in HMXBs were obtained with two methods, namely: disc continuum fitting, and reflection spectroscopy.
 Reflection spectroscopy is complex for soft state spectra (e.g.
 Tomsick et al.
-2018), so here we focus on spin values from disc continuum fitting.
-These are typically very high111The spins attributed through various methods to LMXB BH vary from very low to very high values, sometimes for the same system (e.g.
+2018), so here we focus on spin values from a disc continuum fitting.
+These are typically very high222The spins attributed through various methods to LMXB BH vary from very low to very high values, sometimes for the same system (e.g.
 Reynolds 2021).
@@
 Here, we show that all these BHs may instead form a single population.
-The difference in BH mass arises naturally from the diverse formation environments: all known HMXB-BH are in local group galaxies with relatively high metallicity, so they have strong stellar winds before the core collapse, resulting in lower black hole masses compared to those forming in low metallicity galaxies which are within the sensitivity range of LVK.
+The difference in BH mass arises naturally from the diverse formation environments: all known HMXB-BH systems reside in Local Group galaxies with relatively high metallicity, so they have strong stellar winds before the core collapse.
+This results in lower black hole masses compared to those forming in low metallicity galaxies, which are within the sensitivity range of LVK.
 We demonstrate the difference in expected black hole masses explicitly using a population synthesis model.
 This model also reproduces the spin values in the LVK BHs, for models where high spins are produced only by tidal spin-ups in extremely close binaries.
-The known HMXB-BH systems are in binary systems which are not close enough for tidal spin-up.
-Hence the population synthesis models predict they should not have high spins, which conflicts with the reported high values.
-However, the disc continuum method relies on the emission thermalizing.
-If instead the disc is covered by a warm, optically thick skin, so that the emission is Comptonised, the spin values are all consistent with (Belczynski et al.
+The known HMXB-BH systems are in binary systems that are not close enough for tidal spin-up.
+Hence, the population synthesis models predict they should not have high spins, which conflicts with the reported high values.
+However, the disc continuum method relies on the emission thermalising.
+If, instead, the disc is covered by a warm, optically thick skin, whereby the emission is Comptonised, the spin values are all consistent with (Belczynski et al.
 2021; Zdziarski et al.
 2024a, b; this paper).
-With this interpretation, the LVK BH-BH mergers and the HMXBs can be members of the same binary star population.
+With this interpretation, the LVK BH-BH mergers and the HMXBs can be considered as members of the same binary star population.
 2 BH masses
@@
 2012) that does not produce a lower mass gap between NSs and BHs.
-We use a model in which the pair-instability (upper) mass gap is calculated with non-standard fusion reaction rates and heavy mixing for massive stars that allows for BHs to form up to (Farmer et al.
+We used a model in which the pair-instability (upper) mass gap is calculated with non-standard fusion reaction rates and heavy mixing for massive stars; this allows for BHs to form up to (Farmer et al.
 2020; Costa et al.
 2021).
-Such model was already developed by Belczynski (2020) in the context of the detected at the low metallicity GW190521 LVK merger detected at distance of Gpc Abbott et al.
-(2020)).
-In Figure 1 we show the dependence of the Helium core mass on metallicity for single stars obtained with the stellar wind treatment presented in Belczynski et al.
+This type of model had already been developed by Belczynski (2020) in the context of the detected at the low-metallicity GW190521 LVK merger, at distance of Gpc (Abbott et al.
+2020).
+In Fig.
+1, we show the dependence of the helium core mass on metallicity for single stars obtained with the stellar wind treatment presented in Belczynski et al.
 (2010a).
@@
 The lower the metallicity, the lower the opacity in stellar atmospheres, and the lower the wind mass loss, which produces progressively higher mass BHs with decreasing metallicity.
-Cyg X-1 is located in the Milky Way disc that hosts stars with about solar metallicity (, Asplund et al.
-(2009)).
-For such metallicity, we obtain a maximum BH mass of , which is consistent with the highest mass stellar-origin BH known in the Milky Way for the same Cyg X-1 ( Miller-Jones et al.
-(2021)).
+Cyg X-1 is located in the Milky Way disc, which hosts stars that have metallicity values that are close to solar (, Asplund et al.
+2009).
+For such values, we obtain a maximum BH mass of , which is consistent with the highest mass stellar-origin BH known in the Milky Way for the same Cyg X-1 (, Miller-Jones et al.
+2021).
 LMC X-1 is located in the Large Magellanic Cloud () and M33 X-7 in M33 galaxy ().
 In this metallicity range, BHs are not expected to be more massive than .
-The LVK BH-BH mergers have already been detected (O3) to significant redshifts of .
-The detected BH-BH mergers may have formed even at much larger redshifts due to non-zero delay time between star formation and BH-BH mergers (Dominik et al.
+LVK BH-BH mergers have already been detected (O3) to significant redshifts of .
+The detected BH-BH mergers may have formed even at much larger redshifts due to non-zero delay times between star formation and BH-BH mergers (Dominik et al.
 2012; Fishbach & Kalogera 2021).
-Therefore, galaxies and stars that can produce BH-BH mergers cover a wide range of metallicity.
-It is even claimed that Population III (metal-free) stars () can contribute to LVK detections (Kinugawa et al.
+Therefore, galaxies and stars that can produce BH-BH mergers cover a wide range of metallicities.
+It is even claimed that Population III (metal-free) stars () may contribute to LVK detections (Kinugawa et al.
 2014).
@@
 The recent discovery of a BH in the binary Gaia BH3 (Gaia Collaboration et al.
-2024) strongly supports the scenario according to which metal-poor massive stars are progenitors of the high-mass BHs observed by LVK (see, e.g.
+2024) strongly supports a scenario in which metal-poor massive stars are progenitors of the high-mass BHs observed by LVK (see e.g.
 Olejak et al.
@@
 Population synthesis models are also used to predict the BH spin distribution.
-In most cases, they produce low-spinning BHs () from non-interacting stars and low- to rapidly-spinning BHs () for Wolf-Rayet (WR) stars that were subject to tidal spin-up in close binaries (Belczynski et al.
+In most cases, they produce low-spinning BHs () from non-interacting stars and slow-to-rapidly-spinning BHs () for Wolf-Rayet (WR) stars that were subject to tidal spin-up in close binaries (Belczynski et al.
 2020b; Bavera et al.
@@
 Such models can explain the low and high effective spins of LVK BH-BH mergers (Olejak & Belczynski 2021).
-The low BH spins obtained by population synthesis models are the consequence of the assumed efficient angular momentum transport from stellar cores to envelopes, an effect that has been now well confirmed by observations (Aerts et al.
+The low BH spins obtained by population synthesis models are the consequence of the assumed efficient angular momentum transport from stellar cores to envelopes, an effect that has been now aptly confirmed by observations (Aerts et al.
 2019; Langer 2012) consistent with the action of the magnetic Tayler (1973) instability (Spruit 2002a; Fuller et al.
@@
 2024).
-The heuristic Spruit (2002a) version of the mechanism used in population synthesis codes has been recently confirmed by numerical calculations of Petitdemange et al.
+The heuristic Spruit (2002a) version of the mechanism used in population synthesis codes was recently confirmed by numerical calculations of Petitdemange et al.
 (2024).
-Therefore unless their progenitors in HMXBs belong to a special category of stars with uncoupled cores and envelopes, BHs in these systems should be formed with low spins ().
-The known HMXB orbital periods are too long for tidal interactions to spin up the stars before they form BHs (the size of the present BH’s companion would not fit into a days orbit of the tidal locking regime; see e.g.
+Therefore, unless their progenitors in HMXBs belong to a special category of stars with uncoupled cores and envelopes, the BHs in these systems would have been formed with low spins ().
+The known HMXB orbital periods are too long for tidal interactions to spin up the stars before they form BHs (the size of the present BH companion would not fit into a days orbit of the tidal locking regime; see e.g.
 Belczynski et al.
-2020a) so if their spins are really large they must have have gained angular momentum by accretion.
-To spin up a BH from to one has roughly to double its mass (Bardeen 1970).
-For Cyg X-1 this would require accreting in less than Myr (Miller-Jones et al.
-2021), i.e an accretion rate , or higher by factor of 100 if the accretion lifetime of Cyg X-1 is only a few tens of thousands of years Miller-Jones et al.
+2020a); thus, if their spins are really large, they must have have gained angular momentum by accretion.
+To spin-up a BH from to it is necessary to roughly to double its mass (Bardeen 1970).
+For Cyg X-1, this would require accreting in less than Myr (Miller-Jones et al.
+2021), namely, with an accretion rate of ; this rate would be higher by factor of 100 if the accretion lifetime of Cyg X-1 was only a few tens of thousands of years (Miller-Jones et al.
 2021).
-This gives an accretion rate of where is the Eddington accretion rate with an assumed efficiency of .
+This gives an accretion rate of , where is the Eddington accretion rate with an assumed efficiency of .
 Contrary to the still expressed opinion (see e.g.
 Miller-Jones et al.
-2021), the Eddington rate is not a maximum theoretical limit (Begelman 1979) and black holes can accrete from discs at rates up to at least a few tens of (Kitaki et al.
+2021), the Eddington rate is not a maximum theoretical limit (Begelman 1979) and black holes can accrete from discs at rates of up to (at least) a few tens of (Kitaki et al.
 2021; Yoshioka et al.
 2022, but see Hu et al.
-2022), so that, an accretion rate in Cyg X-1 is not, from this point of view, extravagant.
+2022); thus, an accretion rate in Cyg X-1 would not be, from this point of view, extravagant.
 The problem is to find an evolutionary path that would take a high-mass binary through such a super-Eddington phase and produce the observed Cyg X-1, whose O supergiant donor star has yet to fill its Roche lobe.
-This is unlikely to happen for accretion from wind, even when it is L1-focused (Wiktorowicz et al.
+This is unlikely to happen with respect to accretion from wind, even when it is L1-focused (Wiktorowicz et al.
 2021).
-Therefore, according to the best stellar physics available, BH spins in the known HMXBs should be low, with – consistent with LVK BH spins (Fishbach & Kalogera 2022) and estimates of BH spin in HMXBs by Belczynski et al.
+Therefore, according to the best stellar physics available BH spins in the known HMXBs should be low, with – consistent with LVK BH spins (Fishbach & Kalogera 2022) and estimates of BH spin in HMXBs by Belczynski et al.
 (2021); Zdziarski et al.
-(2024a, b), but clearly incompatible with the generally accepted high spin values in these systems (Miller & Miller 2015).
+(2024a, b); however, this is clearly incompatible with the generally accepted high spin values in these systems (Miller & Miller 2015).
 These high spin values have been obtained by assuming the presence of a standard disc in the system (see Sect.
 7 for the standard-disc definition), which can be justified only if the system is in a soft spectral state.
-In the following we will argue that since Cyg X-1 has never been observed in such a state, the high black hole spin from disc continuum fitting is not a firm requirement.
+In the following, we argue that since Cyg X-1 has never been observed in such a state, the high black hole spin from disc continuum fitting is not a firm requirement.
 4 BH spins from disc spectral fitting: HMXBs
@@
 This gives a total spectrum, which is a sum of all radii of blackbody spectra of different temperatures.
-The emissivity peaks around , so the peak temperature is K, i.e.
-is a factor higher for the same disc luminosity at high spin than at low spin.
-In all X-ray binaries, one observes cycles during which the brightness and spectral hardness form a characteristic “tortoise-head” pattern.
+The emissivity peaks around , so the peak temperature is K, namely, it is a factor higher for the same disc luminosity at high spin than at low spin.
+In all X-ray binaries, we can observe cycles during which the brightness and spectral hardness form a characteristic ‘tortoise-head’ pattern.
 In general, the spectrum contains both thermal disc emission, together with a power law tail extending to much higher energies, clearly showing that there is an additional non-disc component to the accretion flow.
-Most transients show a phase in this cycle where the spectrum is sufficiently soft to be termed disc dominated, where the assumption of accretion disc assumption is likely to be reliable.
+Most transients show a phase in this cycle where the spectrum is sufficiently soft to be termed ‘disc-dominated’, where the accretion disc assumption is likely to be reliable.
 The problem is that not all X-ray sources complete the full cycle.
-As we will show below, this is the case for Cyg X-1.
+As we show below, this is the case for Cyg X-1.
 Indeed, there is clear evidence that in Cyg X-1, the accretion flow has not formed a standard blackbody disc.
-The blue points in Figure 2 show the hardness-intensity diagram of the LMXB GX339-4.
-This source picks out a well-defined track, transitioning from low/hard (lower right) to disc-dominated (upper left, hardness ratio below ), with fast transitions (horizontal) between the two states.
+The blue points in Fig.
+2 show the hardness-intensity diagram of the LMXB GX339-4.
+This source picks out a well-defined track, transitioning from low and hard (bottom-right) to disc-dominated (upper left, hardness ratio below ), with fast transitions (horizontal) between the two states.
 The red points compare data from Cyg X-1 (both data taken from Belloni 2010).
 The mass and distance of GX339-4 are not well known, so we have scaled the relative count rate (brightness) between the two systems to match the lower transition luminosity.
-We can see how GX339-4 transitions from low/hard state (hardness ratio ; the ratio is defined as the ratio of counts in the energy bands and keV) to high/soft state (hardness ratio ).
+We can see how GX339-4 transitions from low/hard state (hardness ratio ; the ratio is defined as the ratio of counts in the energy bands and keV) to high/soft state (hardness ratio of ).
 In contrast, Cyg X-1 never reaches as far to the left as seen in the GX339-4 disc-dominated states.
@@
 2009; Walton et al.
-2016), making it stuck in the more complex intermediate states where there is no reason to believe that the accretion flow forms a standard accretion disc.
-Instead, the soft component in intermediate state spectra is often better fit by optically thick, warm (few keV) Comptonization (see, for e.g.
-transition spectra from the LMXB-BH MAXI J1820+070 in Fig.
+2016), keeping it stuck in the more complex intermediate states where there is no reason to believe that the accretion flow forms a standard accretion disc.
+Instead, the soft component in intermediate state spectra is often better fit by optically thick, warm (few keV) Comptonisation (see e.g.
+transition spectra from LMXB-BH MAXI J1820+070 in Fig.
 2 of Kawamura et al.
@@
 Instead, the strong soft component need not be emitted by a simple blackbody disc, where the energy is dissipated mainly on the midplane but can have heating further up into the photosphere.
-This forms a warm layer on the top of the disc, making it appear hotter and giving higher spin when fit with standard disc models.
+This forms a warm layer on the top of the disc, making it appear hotter and giving a higher spin when fit with standard disc models.
 Such a warm layer on the disc is sometimes required in the black hole binaries (e.g.
-in very high/intermediate type spectra: Kubota et al.
+in very high and intermediate type spectra: Kubota et al.
 2001; Remillard & McClintock 2006) and is generic in active galactic nuclei (Porquet et al.
@@
 2018).
-On the other hand it is well known that the simplistic assumption of dissipation concentrated at the disc midplane (usually in the form of dissipation proportional to pressure) is in contradiction with both observations (see e.g., Hubeny & Long 2021) and numerical simulations (Blaes 2014).
+On the other hand, it is well known that the simplistic assumption of dissipation that is concentrated at the disc midplane (usually in the form of dissipation proportional to pressure) is in contradiction with both observations (see e.g.
+Hubeny & Long 2021) and numerical simulations (Blaes 2014).
 Thus, there is no physical reason to assume that a standard accretion disc emits the observed X-ray spectrum of Cyg X-1.
 In fact, there are good reasons to think that such an assumption is inappropriate for this system.
-There is no “canonical” model describing the inner disc structure in X-ray binaries – no model correctly describes the “tortoise-head” cycle.
-Hence, one is allowed to test various options that would reproduce the observed spectrum and check if the deduced spin value is model-dependent, and if it is, which model corresponds best to the value obtained from stellar evolution.
+There is no ‘canonical’ model describing the inner disc structure in X-ray binaries – no model correctly describes the ‘tortoise-head’ cycle.
+Hence, we can test various options that would reproduce the observed spectrum and check whether the deduced spin value is model-dependent; if it is, we consider which model corresponds best to the value obtained from stellar evolution.
 A similar approach was adopted by Zdziarski et al.
@@
 5 Cyg X-1 spin from the soft component
-We confront two models of Cyg X-1 emission.
-We use data from the softest state ever seen (observed by Suzaku in 2013; magenta star in Fig.
-2), when the soft component is at its most dominant, and so uncertainties from modelling the high energy coronal tail are smallest (observation B4 from Kawano et al.
+We confronted two models of Cyg X-1 emission.
+We used data from the softest state ever seen (observed by Suzaku in 2013; magenta star in Fig.
+2), when the soft component is at its most dominant, and, thus, uncertainties from modelling the high energy coronal tail are smallest (observation B4 from Kawano et al.
 2017, see this paper for details of the data).
 We fit these data using the xspec spectral fitting package and all the models below include absorption along the line of sight using tbabs with column density as a free parameter (see Appendix A for the details).
-We show results from a series of model fits in Figure 3, where the original (absorbed) X-ray Suzaku data is grey, with the reconstructed (corrected for absorption) spectrum shown with orange points.
-We assume here the standard mass and distance, and fix the disc inclination to the binary inclination of .
+We show results from a series of model fits in Fig.
+3, where the original (absorbed) X-ray Suzaku data is grey, with the reconstructed (corrected for absorption) spectrum shown with orange points.
+We assume here the standard mass and distance, and we fix the disc inclination to the binary inclination of .
 We first fit the standard disc model to these data.
-The results are shown in the left panel of Figure 3.
+The results are shown in the left panel of Fig.
+3.
 This includes all special and general relativistic ray tracing effects, as well as allowing a correction to the blackbody temperatures to account for the incomplete thermalisation expected from full models of the vertical disc structure (colour temperature correction, fixed to ).
-We allow a fraction of these disc spectra to be Compton upscattered to form the power law tail, modelled using the standard formalism, which assumes the electron energy is typically higher than keV.
-We recover the usual, very high black hole spin () seen in all standard-disc fits to the soft component (Zhao et al.
+We allowed for a fraction of these disc spectra to be Compton upscattered to form the power law tail, modelled using the standard formalism, which assumes the electron energy is typically higher than keV.
+We recovered the usual, very high black hole spin () seen in all standard-disc fits to the soft component (Zhao et al.
 2021a; Zdziarski et al.
 2024a, b).
-The middle panel of Figure 3 shows what happens when we additionally allow a warm skin on the top of the disc.
-We model this using another Compton upscattering model, which allows the electron temperature to be a free parameter, as well as the scattered fraction (set by the optical depth of the skin).
-This is formally a much better fit to the data, even accounting for the additional free parameters: optical depth and temperature of the warm disc skin.
+The middle panel of Fig.
+3 shows what happens when we additionally allow a warm skin on the top of the disc.
+We model this using another Compton upscattering model, which allows the electron temperature to remain as a free parameter, as well as the scattered fraction (set by the optical depth of the skin).
+This is formally a much better fit to the data, even accounting for the two additional free parameters: optical depth and temperature of the warm disc skin.
 A simple F-test gives a probability of the improvement being random as .
@@
 The BH spin here is extremely poorly constrained as the soft component shape in this model is set more by the electron temperature of the skin than by the temperature of the disc at the ISCO.
-We illustrate this in the right panel of Figure 3, where we show a fit with almost the same but for , as predicted by standard stellar evolution.
-This demonstrates that formal solutions for BH spin from spectra with their reported extremely small error bars or very definitive values (e.g., at level (Zhao et al.
-2021b)) are model-dependent and are not the only solution (see also Zdziarski et al.
+We illustrate this in the right panel of Fig.
+3, where we show a fit with almost the same but for , as predicted by standard stellar evolution.
+This demonstrates that formal solutions for BH spin from spectra with their reported extremely small error bars or very definitive values (e.g.
+at level, Zhao et al.
+2021b) are model-dependent and are not the only solution (see also Zdziarski et al.
 2024a, b who arrive at the same conclusion).
-5.1 Can a disc form in Cyg X-1 for low BH spin?
-Modelling of the disc continuum spectra assumes that an accretion disc can form around the BH for any value of the BH spin and/or any orientation of the accretion disc with respect to the spin of the BH.
+5.1 Considering whether a disc can form in Cyg X-1 for a low BH spin
+The modelling of the disc continuum spectra assumes that an accretion disc can form around the BH for any value of the BH spin and/or any orientation of the accretion disc with respect to the spin of the BH.
 In a binary system, a disc will form only if the circularisation radius of the transferred matter is larger than the radius of the accreting body.
 In other words, a disc can form if the angular momentum of the accreting matter is large enough to allow it to circle around the accretor.
-For wind (Bondi-Hoyle; Bondi 1952) accretion the condition is (Frank et al.
-2002)
-where is the mass of the donor star, the orbital separation, the wind velocity and the escape velocity from the donor.
-For the parameters of Cyg X-1, when , cm but cm, so that e.g.
-for no disc can form.
+For wind accretion (Bondi-Hoyle; Bondi 1952), the condition is (Frank et al.
+2002):
+where is the mass of the donor star, the orbital separation, the wind velocity, and the escape velocity from the donor.
+For the parameters of Cyg X-1, when , cm but cm, so that, for instance, for no disc can form.
 Sen et al.
@@
 A high Roche lobe filling factor has been found for LMC X-1 (Orosz et al.
-2009) and the O donor in M33 X-7 effectively fills its Roche lobe (although not in a standard way, see Ramachandran et al.
+2009) and the O donor in M33 X-7 effectively fills its Roche lobe (although not in a standard way; see Ramachandran et al.
 2022).
@@
 The high Roche lobe filling factor of the O star enables the formation of a focused wind with sufficient angular momentum to make a prograde accretion disc.
-In such case the circularisation radius for the bulk fo the accretion flow is of the order of the solar radius (see, e.g., Frank et al.
+In such cases, the circularisation radius for the bulk of the accretion flow is of the order of the solar radius (see e.g.
+Frank et al.
 2002), orders of magnitude larger than the ISCO radius.
-The above arguments, together with our modelling of the disc continuum spectra imply that the observations of Cyg X-1 are consistent with the statement that its BH can also be slowly spinning.
+The above arguments, together with our modelling of the disc continuum spectra, imply that the observations of Cyg X-1 are consistent with the notion that its BH could also be slowly spinning.
 6 M33 X-7 BH spin from the soft component
@@
 With this new spin value, the M33 X-7 BH is no longer in the very-rapidly rotating class, to which Cyg X-1 with and LMC X-1 with belong.
-Nevertheless, since we have concluded above that the spin determination in Cyg X-1 is strongly model-dependent and Zdziarski et al.
-(2023, 2024b) arrived to same conclusion regarding LMC X-1, we thought it useful to check if the same applies to M33 X-7.
-Similarly to these sources, when a warm-skin Comptonizing model is used, the black hole spin is completely unconstrained, with acceptable values ranging from -1 to at confidence level (see Appendix B).
-Below, we will discuss the possible reason for this common property of the only three HMBXs that are potential progenitors of BHBH binary systems.
+Nevertheless, since we have concluded that the spin determination in Cyg X-1 is strongly model-dependent, supported by the findings of Zdziarski et al.
+(2023, 2024b), who arrived to same conclusion regarding LMC X-1, we thought it useful to check if the same applies to M33 X-7.
+Similarly to these sources, when a warm-skin Comptonising model is used, the black hole spin is completely unconstrained, with acceptable values ranging from -1 to at confidence level (see Appendix B).
+Below, we discuss the possible reason for this common property of the only three HMBXs that are potential progenitors of BHBH binary systems.
 7 Discussion
-We have shown that the X-ray spectra of Cyg X-1 and M88 X-7 can be fitted with both standard disk and warm Comptonised disc models (see also Zdziarski et al.
+We have shown that the X-ray spectra of Cyg X-1 and M33 X-7 can be fitted with both standard disc and warm Comptonised disc models (see also Zdziarski et al.
 2024a, b for the same conclusion about Cyg X-1 and LMC X-1).
 We have also noticed that even in its softest state, the spectrum of Cyg X-1 is not soft enough to justify a standard-disc model fit, so there are very good reasons to consider alternatives.
-In order to avoid ambiguities and misunderstandings, one has to define precisely what is understood by fitting spectra with a “standard disc” model.
-This model is describing an axially symmetric, stationary, geometrically thin Keplerian accretion flow.
-It is assumed that the angular momentum transport is provided by a “viscous” (in astrophysical discs, the viscosity must be anomalous) torque and that the local dissipation of the same torque is the source of the energy radiated away from the disc surface.
-Then, from mass and angular momentum conservation equations for the vertically averaged disc parameters, assuming thermal equilibrium, one obtains
+In order to avoid ambiguities and misunderstandings, one has to define precisely what is understood by fitting spectra with a ‘standard disc’ model.
+This model describes an axially symmetric, stationary, geometrically thin Keplerian accretion flow.
+It is assumed that the angular momentum transport is provided by a ‘viscous’ (noting that in astrophysical discs, the viscosity must be ’anomalous’) torque and that the local dissipation of the same torque is the source of the energy radiated away from the disc surface.
+Then, from the mass and angular momentum conservation equations for the vertically averaged disc parameters, assuming thermal equilibrium, we obtain
 where is the effective temperature and contains information about the boundary conditions.
-This is the equation that provides the flux, or effective temperature profile for a Keplerian stationary disc.
+This is the equation that provides the flux – or effective temperature profile, – for a Keplerian stationary disc.
 It does not say that the disc emits radiation like a blackbody.
-It contains no information about the vertical energy transport in the disc and is independent of the angular momentum and dissipation mechanism it only assumes the existence of such mechanism (it is independent of the value of the “viscosity parameter , it does not even assume the existence of such a parameter).
-Eq.
-(3) is the basis of the “standard” disc X-ray spectrum model.
-There, one assumes that the colour temperature is close to the effective temperature.
-i.e., that the disc, locally radiates approximately like a black body.
+It contains no information about the vertical energy transport in the disc and is independent of the angular momentum and dissipation mechanism; in fact, it only assumes the existence of such mechanism (independently of the value of the viscosity parameter, , and does not even assume the existence of such a parameter).
+Equation (3) is the basis of the ‘standard’ disc X-ray spectrum model.
+There, it is assumed that the colour temperature is close to the effective temperature; namely, the disc, locally radiates approximately as a blackbody would.
 Since the accretor is a BH, one adds to the model general-relativistic effects, including radiation back-falling onto the disc, and appropriate inner boundary conditions (usually zero torque at ISCO).
-Since the observed spectra are not strictly a sum of black-bodies, a disc annulus emission is represented by a colour-corrected (or diluted) black-body
-where is the intensity, is the Planck spectrum, the Planck constant, the radiation frequency and is a colour temperature correction factor representing the extent to which electron scattering through the disc vertical structure is expected to affect the emergent spectrum.
+Since the observed spectra are not strictly a sum of blackbodies, the disc annulus emission is represented by a colour-corrected (or diluted) blackbody:
+where is the intensity, is the Planck spectrum, the Planck constant, the radiation frequency, and is a colour temperature correction factor representing the extent to which electron scattering through the disc vertical structure is expected to affect the emergent spectrum.
 This can be calculated by solving the disc vertical radiative transfer equations using a code such as TLUSTY (see Davis & El-Abd 2019; Davis et al.
 2005 and references therein).
-However, contrary to to stellar atmospheres, accretion discs are not in radiative equilibrium, i.e.
-the divergence of the radiative flux is not equal to zero, but to the (viscous) heating rate per unit volume.
-Hence the calculations requires the viscous–dissipation vertical structure as input.
-In the -model of Shakura & Sunyaev (1973), one assumes that this dissipation profile is identical to that of the total pressure, but this is radiation pressure unstable at the observed disc luminosities.
-Instead, for the disc calculations, the dissipation is assumed to follow the density profile in TLUSTY Done & Davis (2008).
-Both these prescriptions have maximal dissipation at the mid-plane, whereas MRI simulations often show maximum dissipation closer to the disc surface (see, e.g., Blaes 2014).
-Some versions of TLUSTY (Hubeny & Hubeny 1998) use a step-wise power-law dependence of viscosity that allows a chosen fraction of the dissipation to occur in the surface layers.
-Such dissipation profile is needed to reproduce disc spectra of cataclysmic variables Hubeny & Long (2021) and when applied to BH X-ray emitting disc, such dissipation stratification produces disc coronas (Davis & El-Abd 2019) as suggested by Svensson & Zdziarski (1994).
-What is, or rather what are the real, physical dissipation profiles is still unknown since numerical simulations cannot yet produce realistic vertical structures of geometrically thin accretion discs.
-That is why, disc models with warm surface layers are as physically legitimate as the presumed “standard” model, especially when there is strong evidence as in Cyg X-1 that the structure is not a standard disc.
-One should also point out that the -disc model cannot explain the brightness–hardness ratio tortoise-head hysteresis observed in transient X-ray sources.
+However, contrary to stellar atmospheres, accretion discs are not in radiative equilibrium, namely, the divergence of the radiative flux is not equal to zero, but to the (viscous) heating rate per unit volume.
+Hence, the calculations requires the viscous-dissipation vertical structure as input.
+In a simplistic application of the -model of Shakura & Sunyaev (1973), it is assumed that this dissipation profile is identical to that of the total pressure.
+Instead, for the disc emission calculations, the dissipation is assumed to follow the density profile in TLUSTY (Done & Davis 2008).
+Both these prescriptions have maximal dissipation at the mid-plane, whereas MRI simulations often show maximum dissipation closer to the disc surface (see e.g.
+Blaes 2014).
+Some versions of TLUSTY (Hubeny & Hubeny 1998) use a step-wise power-law dependence of viscosity that allows for a chosen fraction of the dissipation to occur in the surface layers.
+Such a dissipation profile is needed to reproduce disc spectra of cataclysmic variables Hubeny & Long (2021) and when applied to BH X-ray emitting disc, such dissipation stratification produces disc coronas (Davis & El-Abd 2019), as suggested by Svensson & Zdziarski (1994).
+In this regard, the real, physical dissipation profiles are still unknown since numerical simulations cannot yet produce realistic vertical structures of geometrically thin accretion discs.
+That is why disc models with warm surface layers are as physically legitimate as the presumed ‘standard’ model, especially when there is strong evidence as in Cyg X-1 that the structure is not a standard disc.
+It is also useful to point out that the -disc model cannot explain the brightness-hardness ratio tortoise-head hysteresis observed in transient X-ray sources.
 The physical reason for this variability is still unknown.
@@
 2007; Basak & Zdziarski 2016; Basak et al.
-2017) and models (e.g., Dubus et al.
+2017) and models (e.g.
+Dubus et al.
 2001) suggests that accretion discs in BH X-ray binaries are truncated, but the physical mechanism of this excision is still a mystery.
-It does not mean that the -disc model on which the “standard” spectral model is based, is to be rejected.
+This does not mean that the -disc model on which the ‘standard’ spectral model is based is to be rejected.
 This model has found many successful applications.
-The soft state spectra in transients show peak disk temperature and total luminosity changing together with , indicating a constant inner radius for the disc and constant Done et al.
-(2007).
-Even the radial dependence of the disc temperature shows the predicted behaviour in eclipsing cataclysmic variables in their disk (outbursting) states (Horne & Cook 1985), as well as showing the very different radial profile expected from non-equilibrium discs in quiescent dwarf-novae; Lasota 2001.
-None of the known HMBH show a convincing soft state relation which characterises a standard disc (see Sect.
+The soft state spectra in transients show peak disk temperature and total luminosity changing together with , indicating a constant inner radius for the disc and constant (Done et al.
+2007).
+The radial dependence of the disc temperature shows the predicted behaviour in eclipsing cataclysmic variables in their disk (outbursting) states (Horne & Cook 1985), as well as showing the very different radial profile expected from non-equilibrium discs in quiescent dwarf-novae; (Lasota 2001).
+None of the known HMBHs exhibit a convincing soft state relation that characterises a standard disc (see Sect.
 4).
-Cyg X-1 is never in a soft state, while in LMC X-1, Gierliński & Done (2004a) discovered that the disc luminosity follows instead a track.
+Cyg X-1 is never in a soft state, while in LMC X-1 Gierliński & Done (2004a) discovered that the disc luminosity follows instead a track.
 Zdziarski et al.
-(2024a) found that such a anti-correlation between luminosity and temperature can be naturally explained as an effect of the warm Comptonization layer on top of a standard disc.
-The luminosity variations in M88 X-7 are not large enough to determine whether a relation is present 222Liu et al.
+(2024a) found that such a anti-correlation between luminosity and temperature can be naturally explained as an effect of the warm Comptonisation layer on top of a standard disc.
+The luminosity variations in M33 X-7 are not large enough to determine whether a relation is present 333Liu et al.
 (2008) present a variability in counts (not in counts per second) which reflects mostly the variability of the exposure time..
-The three BHHMBXs, Cyg X-1, LMC X-1 and M33 X-7, whose spectra are well fitted by a model of a disc covered with warm comptonising layer have one feature in common: the accrete not only from focused stream but also capture some matter from the strong wind of the BH companion.
+The three BHHMBXs, Cyg X-1, LMC X-1, and M33 X-7, whose spectra are well fitted by a model of a disc covered with warm comptonising layer have one feature in common: they accrete not only from a focused stream, but also capture some matter from the strong wind of the BH companion.
 It is natural to speculate that this is the source of the warm layer, making it a ubiquitous feature of the HMBH spectra.
@@
 The difference between HMXB and LVK BH masses has already been noticed after the first gravitational wave detection and attributed to metallicity effect (Abbott et al.
-2016) as anticipated by Belczynski et al.
+2016), as anticipated by Belczynski et al.
 (2010b).
@@
 1).
-While assessing the problem of BH mass difference between HMXB and gravitational–wave sources is rather straightforward, addressing the apparent conflicting spin values is more complicated.
+Although assessing the problem of BH mass differences between HMXB and gravitational–wave sources is rather straightforward, the task of addressing the apparent conflicting spin values is more complicated.
 At face value, the spin values attributed to the BH in the HMXBs that are massive enough to form LVK sources (Cyg X- 1, LMC X-1, M33 X-7) seem to require two distinct populations: slowly spinning (LVK) and rapidly spinning (HMXBs) BHs.
-Since the majority of LVK BHs are slowly spinning (), we argue, under the assumption that they are stellar-origin BHs (i.e., are not primordial BHs: Hawking 1971, but see Mroz et al.
-2024), that it must indicate that massive stars are subject to efficient angular momentum transport (Spruit 2002b; Fuller et al.
+Since the majority of LVK BHs are slowly spinning (), we argue, under the assumption that they are stellar-origin BHs (i.e.
+they are not primordial BHs: Hawking 1971, but see Mróz et al.
+2024), that this must indicate that massive stars are subject to efficient angular momentum transport (Spruit 2002b; Fuller et al.
 2019; Eggenberger et al.
@@
 2023).
-These moderately/rapidly spinning BHs can also be explained in the framework of efficient angular momentum transport.
+These moderately and rapidly spinning BHs can also be explained in the framework of efficient angular momentum transport.
 Although the star rotation slows down during its evolution towards the formation of a BH, the tidal spin-up of a WR star (BH progenitor) in close BH+WR binary systems (BH-BH progenitors) can easily result in a moderately or rapidly spinning BH (Olejak & Belczynski 2021).
-In dense clusters, dynamical interactions may lead to the formation of BH-BH mergers with components that are second or even third-generation BHs (formed by earlier BH-BH mergers) that naturally have significant spins (Rodriguez et al.
+In dense clusters, dynamical interactions may lead to the formation of BH-BH mergers with components that are second- or even third-generation BHs (formed by earlier BH-BH mergers) that naturally exhibit significant spins (Rodriguez et al.
 2019).
-Under the natural assumption that BHs in HMXBs are formed directly from massive stars and within the paradigm of efficient angular momentum transport, these BHs also must have low spins (; e.g., Belczynski et al.
+Under the natural assumption that BHs in HMXBs are formed directly from massive stars and within the paradigm of efficient angular momentum transport, these BHs also must have low spins (; e.g.
+Belczynski et al.
 2020b).
-We have demonstrated that this conclusion is fully consistent with BH spin estimate based on fitting the disc spectra in HMXBs with models in which the disc is covered by a warm Comptonized layer.
-In particular, we show that with this model, the X-ray spectra of the softest, but still pretty hard, state of Cyg X-1 is consistent with a slowly spinning BH as confirmed by Zdziarski et al.
+We have demonstrated that this conclusion is fully consistent with BH spin estimate based on a fitting of the disc spectra in HMXBs with models where the disc is covered by a warm Comptonised layer.
+In particular, we show that with this model, the X-ray spectra of the softest (but still pretty hard) state of Cyg X-1 is consistent with a slowly spinning BH, as confirmed by Zdziarski et al.
 (2024b) for another set of data.
-The same is true of X-7 in M33 whose BH spin is completely unconstrained when its soft X-ray spectra are fitted with the warm–skin model.
-The same warm–skin solutions also give a low spin for LMC X-1 (Zdziarski et al.
+The same is true of X-7 in M33, whose BH spin is completely unconstrained when its soft X-ray spectra are fitted with the warm-skin model.
+The same warm-skin solutions also give a low spin for LMC X-1 (Zdziarski et al.
 2024a.
-We have shown that the warm–skin solution is as physically plausible as the so–called standard solution, especially in the case of systems that are not in a generic soft X-ray spectral state.
+We have shown that the warm-skin solution is as physically plausible as the so–called standard solution, especially in the case of systems that are not in a generic soft X-ray spectral state.
 We speculate that this warm skin is ubiquitous in the HMXB-BH due to the presence of additional accretion from lower angular momentum material in the stellar wind as well as Roche Lobe overflow.
@@
 - El Mellah et al. (2019) El Mellah, I., Sander, A. A. C., Sundqvist, J. O., & Keppens, R. 2019, A&A, 622, A189
-- Farmer et al. (2020) Farmer, R., Renzo, M., de Mink, S., Fishbach, M., & Justham, S. 2020, arXiv e-prints, arXiv:2006.06678
+- Farmer et al. (2020) Farmer, R., Renzo, M., de Mink, S. E., Fishbach, M., & Justham, S. 2020, ApJ, 902, L36
 - Fishbach & Kalogera (2021) Fishbach, M. & Kalogera, V. 2021, ApJ, 914, L30
@@
 - Fryer et al. (2012) Fryer, C. L., Belczynski, K., Wiktorowicz, G., et al. 2012, ApJ, 749, 91
-- Fuller et al. (2019) Fuller, J., Piro, A. L., & Jermyn, A. S. 2019, MNRAS[arXiv:1902.08227]
+- Fuller et al. (2019) Fuller, J., Piro, A. L., & Jermyn, A. S. 2019, MNRAS, 485, 3661
 - Gaia Collaboration et al. (2024) Gaia Collaboration, Panuzzo, P., Mazeh, T., et al. 2024, arXiv e-prints, arXiv:2404.10486
@@
 - Miller-Jones et al. (2021) Miller-Jones, J. C. A., Bahramian, A., Orosz, J. A., et al. 2021, Science, 371, 1046
-- Mroz et al. (2024) Mroz, P., Udalski, A., Szymanski, M. K., et al. 2024, arXiv e-prints, arXiv:2403.02386
+- Mróz et al. (2024) Mróz, P., Udalski, A., Szymański, M. K., et al. 2024, Nature, 632, 749
 - Novikov & Thorne (1973) Novikov, I. D. & Thorne, K. S. 1973, in Black Holes (Les Astres Occlus), 343–450
@@
 Appendix A Fitting Cyg X-1 spectra
-We fit Cyg X-1 using two models: one considering a standard disc, and another where the disc is covered by some warm optically thick material (model a and b respectively).
+We fit Cyg X-1 using two models: 1) considering a standard disc and 2) where the disc is covered by some warm optically thick material (models a and b, respectively).
 In both cases we model the spectra using xspec v.12.13.0 (Arnaud 1996).
@@
 2005).
-This considers multi-colour black-body emission originating from a disc structure, assuming Novikov & Thorne (1973) emissivity.
+This considers multi-colour blackbody emission originating from a disc structure, assuming Novikov & Thorne (1973) emissivity.
 All relativistic effects are taken into account, including ray-tracing from the disc to the observer.
@@
 2009).
-This takes an input seed spectrum (in this case kerrbb) and scatters a fraction of the seed photons into a power-law component (hence emulating a Comptonized spectrum).
+This takes an input seed spectrum (in this case kerrbb) and scatters a fraction of the seed photons into a power-law component (hence emulating a Comptonised spectrum).
 In xspec syntax, the model is simpl*kerrbb.
@@
 1996, 2020).
-This takes an input seed spectrum and Compton scatters to higher energies, similar to simpl, however unlike simpl thcomp takes the electron temperature of the Comptonizing plasma as an input argument, setting the high energy turn over.
+This takes an input seed spectrum and Compton scatters to higher energies, similar to simpl, however unlike simpl thcomp takes the electron temperature of the Comptonising plasma as an input argument, setting the high energy turn over.
 Additionally, for this optically thick skin we assume that all input seed photons are Compton scattered, unlike the case for the high-energy tail where only a fraction are scattered.
@@
 We note that in Fig.
-3 we show the spectrum corrected for absorption in colour, and the original uncorrected spectrum in grey.
-As the data originate from two different instruments, XIS (soft-xrays) and HXD (hard-xrays), though both onboard Suzaku, we incldue a cross-calibration constant in all our models.
-Throughout this is fixed to 1.15, i.e increasing the model flux in the hard band by a factor 1.15.
+3 we show the spectrum corrected for absorption in colour and the original uncorrected spectrum in grey.
+As the data originate from two different instruments, XIS (soft-xrays) and HXD (hard-xrays), while both onboard Suzaku, we included a cross-calibration constant in all our models.
+Throughout, this was fixed to 1.15, that is: increasing the model flux in the hard band by a factor 1.15.
 Finally, the total xspec models are: const*tbabs*csimpl*kerrbb for model a, and const*tbabs*csimpl*thcomp*kerrbb for model b.
-Appendix B Re-Visiting M33-X7
-Here we re-visit the black hole spin estimate of M33-X7, performed by Liu et al.
+Appendix B Revisiting M33-X7
+Here we revisit the black hole spin estimate of M33-X7, performed by Liu et al.
 (2008), using the same methodology as for Cyg X-1.
@@
 (2008).
-Initial inspection of the data show strongly disc dominated spectra, with no clear non-thermal high energy tail within the observed energy range.
-Hence we use a slightly simpler model than with Cyg X-1, modelling the continuum as either a pure disc (using kerrbb, Li et al.
+An initial inspection of the data show strongly disc dominated spectra, with no clear non-thermal high energy tail within the observed energy range.
+Hence, we used a slightly simpler model than with Cyg X-1, modelling the continuum as either a pure disc (using kerrbb, Li et al.
 (2005)) or a disc with a warm Comptonising medium above it (using thcomp*kerrbb, Zdziarski et al.
@@
 For completeness we fit the data using both the new and previous mass estimates, to highlight that our conclusions are not sensitive to this change.
-Fig.
-4 shows the four ACIS spectra along with their respective model fits, for the pure disc (left) and warm Comptonised medium (right).
-The data show a change in normalisation between some of the spectra, corresponding to a change in the mass-accretion rate (see e.g the difference between the green/black data/model and the magenta/orange data/model in Fig.
+Figure 4 shows the four ACIS spectra along with their respective model fits, for the pure disc (left) and warm Comptonised medium (right).
+The data show a change in normalisation between some of the spectra, corresponding to a change in the mass-accretion rate (see e.g the difference between the green-black data model and the magenta-orange data model in Fig.
 4).
@@
 The spectra were grouped such that each energy bin contained a minimum of 20 counts, allowing for the use of statistics.
-Fig.
-5 shows the change in as we step the fit through the possible values for black hole spin, with a standard disc on the left and a disc covered by a warm Comptonising material on the right.
+Figure 5 shows the change in as we step the fit through the possible values for black hole spin, with a standard disc on the left and a disc covered by a warm Comptonising material on the right.
 The standard disc clearly gives a strongly constrained value for the black hole spin, and in fact for the old mass estimate we recover the value derived in Liu et al.